OpenStudy (anonymous):

1/6 divided by 3/10 in lowest terms and as a whole or mixed number if possible.. Can someone please help me with this problem. ASAP. Thanks

OpenStudy (anonymous):

1/6 divided by 3/10 in lowest terms and as a whole or mixed number if possible.. Can someone please help me with this problem. ASAP. Thanks

Directrix (directrix):

1/6 divided by 3/10 ----------- 1/6 divided by 3/10 = 1/6 times 10/3 = 10 / 18 = 5 / 9 .